# Multi-Draft Speculative Sampling: Canonical Architectures and Theoretical Limits

### v1

@@ Abstract
We consider multi-draft speculative sampling, where the proposal sequences are sampled independently from different draft models. At each step, a token-level draft selection scheme takes a list of valid tokens as input and produces an output token whose distribution matches that of the target model. Previous works have demonstrated that the optimal scheme (which maximizes the probability of accepting one of the input tokens) can be cast as a solution to a linear program. In this work we show that the optimal scheme can be decomposed into a two-step solution: in the first step an importance sampling (IS) type scheme is used to select one intermediate token; in the second step (single-draft) speculative sampling is applied to generate the output token. For the case of two identical draft models we further 1) establish a necessary and sufficient condition on the distributions of the target and draft models for the acceptance probability to equal one and 2) provide an explicit expression for the optimal acceptance probability. Our theoretical analysis also motives a new class of token-level selection scheme based on weighted importance sampling. Our experimental results demonstrate consistent improvements in the achievable block efficiency and token rates over baseline schemes in a number of scenarios.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has revolutionized the field of natural language processing and deep learning. One of the key factors contributing to the success story of transformers, as opposed to prior recurrent-based architectures (Hochreiter and Schmidhuber, 1997; Chung et al., 2014), is their inherent train-time parallelization due to the attention mechanism. This allows for massive scaling and lead to the development of state-of-the-art Large Language Models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Brown et al., 2020; Chowdhery et al., 2023) which have demonstrated remarkable performance across a wide range of tasks. Despite their parallelizable training, LLM inference is sequential, owing to their auto-regressive nature. This limits their text-generation to one token per one forward pass, which is known to be memory-bound (Shazeer, 2019).
To alleviate the memory-bound nature of auto-regressive decoding of LLMs, speculative decoding (Chen et al., 2023; Leviathan et al., 2023) leverages an arbitrary smaller language model (draft model) that generates multiple candidate tokens in an auto-regressive manner. The LLM (target model) is then used to score all the tokens in the draft in parallel, and the draft tokens are verified through a sequence of token-level rejection sampling which guarantees that the final sequence follows the same distribution as that of the target model. In order for speculative decoding to be beneficial, the combined cost of auto-regressively sampling from the draft model and parallel verification via the target model should be smaller than auto-regressively sampling from the target model. Intuitively, this requires that the draft model distribution resembles that of the target model, which can be measured via the acceptance rate of the speculative decoding process, i.e., the rate at which we accept/reject draft tokens.
A large number of works on speculative decoding (Sun et al., 2024b; Jeon et al., 2024; Miao et al., 2024; Sun et al., 2024a) have emerged recently in an effort to further improve decoding efficiency. The authors in (Sun et al., 2024b) propose SpecTr, a multi-draft extension where the draft model generates $K$ candidate token sequences (which could be sampled in a batch) for each time-step (as opposed to one). The authors consider a token-level selection scheme with the objective of maximizing the probability of accepting some token in the set of available tokens. They demonstrate that this problem can be cast into the framework of optimal transport and solved using a linear program. However due to complexity reasons, the authors instead propose a modified sequential rejection sampling scheme.
@@ 1.1 Main Contributions
  • We revisit the optimal transport framework introduced in  (Sun et al., 2024b) and introduce a canonical architecture — we demonstrate that the optimal acceptance probability can be achieved by a two-step scheme: the first step involves selecting a token from the input set using a type of importance sampling; the second step involves speculative sampling using the selected token and the target distribution.
@@ 2 Background and Related Work
Auto-regressive sampling from LLMs is inherently sequential and memory-bound (Shazeer, 2019). Several approaches have been proposed in the literature to accelerate LLM inference (Shazeer, 2019; Jaszczur et al., 2021; Frantar et al., 2022; Frantar and Alistarh, 2023; Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023; Jeon et al., 2024; Sun et al., 2024b; Miao et al., 2024). Model compression techniques, such as quantization (Frantar et al., 2022; Bondarenko et al., 2024) and sparsification (Jaszczur et al., 2021; Frantar and Alistarh, 2023) have been shown to reduce the overall complexity of LLMs at the expense of some degradation in decoding quality.
@@
For lossless LLM inference acceleration, speculative decoding (Chen et al., 2023; Leviathan et al., 2023; Stern et al., 2018) has emerged as a promising and orthogonal alternative. Earlier works on greedy decoding can draft and predict multiple tokens by augmenting the base LLM (Stern et al., 2018) or aggressive decoding (Ge et al., 2022). However, LLM text-generation often requires sampling with non-zero temperature from the generated logits. To that end, speculative decoding (Chen et al., 2023; Leviathan et al., 2023) was proposed. In speculative decoding, auto-regressive sampling is delegated to a smaller language model (draft model) that generates multiple candidate tokens. The LLM (target model) is then used to score all the tokens in the draft in parallel, and the draft tokens are verified through a sequence of token-level rejection sampling. Speculative decoding guarantees that the final sequence follows the same distribution as that of the target model. The performance of speculative methods highly depends on the choice of the draft model. Zhou et al. (2023) use knowledge distillation (Hinton et al., 2015) to better align the draft and target models which results in higher token acceptance rates.
@@
More recently, the works of Sun et al. (2024b); Miao et al. (2024); Jeon et al. (2024) extend speculative decoding to the multi-draft setting where the draft model(s) generate multiple token sequences per time-step. Specifically, Sun et al. (2024b) formulate the token-level draft selection problem as a discrete optimal transport problem with membership cost and propose SpecTr: a new decoding algorithm that allows for multiple candidates for each token in the draft. A related setting is also studied in Miao et al. (2024); Jeon et al. (2024) where the authors consider a token tree based construction for improving the draft sequences as well as a token-level selection method different form Sun et al. (2024b) . Instead of using a dedicated draft model, Cai et al. (2024) propose augmenting the target model with extra decoding heads that can concurrently draft multiple tokens. The extra heads are fine-tuned using parameter-efficient methods, and can be added to any pre-trained target model. Orthogonally, Sun et al. (2024a) study block-level verification in the single-draft setting as a block-level optimal transport problem. They propose a computationally-efficient algorithm that optimally solves the block-level transport problem, and report speedups over prior token-level verification (Leviathan et al., 2023).
@@ 3 Token-Level Optimal Draft Selection: Theoretical Analysis
We focus on token-level optimal draft selection framework introduced in Sun et al. (2024b). We assume that $\Omega=\{1,2,\ldots,n\}$ denotes the vocabulary of tokens and at a given step, say $t$, ${\mathcal{S}}=\{X_{1},\ldots,X_{K}\}$, denotes the $K$ valid tokens under consideration. Each of these tokens is generated in an i.i.d. fashion from a distribution $p(\cdot)$ determined by the underlying draft modelWe also consider the case when the tokens are generated from different distributions. See Remarks 1 and 3, as well as the experimental results in Section 5. and the context sequence $u^{t}\in\Omega^{t}$ i.e, for each $y\in\Omega$, we have $p(y)={\mathcal{M}}_{s}(y|u^{t})$, where ${\mathcal{M}}_{s}$ denotes the distribution generated by the small (draft) model. In a similar fashion we let $q(\cdot)$ be the distribution over $\Omega$ associated with the large model i.e., $q(y)={\mathcal{M}}_{b}(y|u^{t})$ where ${\mathcal{M}}_{b}$ denotes the distribution generated by the large model. Note that we do not explicitly indicate the sequence $u^{t}$ when discussing $p(\cdot)$ and $q(\cdot)$, as it is fixed and common to both models throughout our analysis.
@@
Problem 1 was studied in Sun et al. (2024b) and shown to be an instance of optimal transport, which can be cast as a linear program. The authors used this framework to establish the optimality of speculative sampling (Chen et al., 2023; Leviathan et al., 2023) in the case of a single draft i.e., $K=1$. For $K>1$ the authors established an information theoretic upper bond on $P^{\star}(\mathrm{acc})$. In this work, we revisit Problem 1 and develop new insights into the structure of the optimal solution. In fact, we establish that the optimal solution in the case of multiple drafts has a natural connection to importance sampling (Tokdar and Kass, 2010). For the case of $K=2$ drafts we exactly characterize $P^{\star}(\mathrm{acc})$ and state necessary and sufficient conditions on $p(\cdot)$ and $q(\cdot)$ for $P^{\star}(\mathrm{acc})$ to equal $1$.
@@
Note that instead of considering the probability over the value of the selected token in (1), one can instead consider the probability of selecting an index $i$ between $\{1,\,\ldots,K\}$ i.e., $\Pr(I=i|X_{1:K}=x_{1:K})$. Such a distribution maps to (1) by simply summing over all indices where $x_{i}=y$. We note that the form in (1) will be more convenient in the sequel. Also note that the classical importance sampling scheme (Tokdar and Kass, 2010) corresponds to the case where $\Pr(I=i|X_{1}^{k}=x_{1:K})\propto q(x_{i})/p(x_{i})$. However the family of schemes in Definition 1 is not restricted to such a choice and we treat $\beta_{y}(x_{1},\ldots,x_{K})$ as free parameters that can be optimized. While an importance weighted sampling scheme may not lead be a valid TLSR, as explained next, it is a key building block in a canonical architecture for token selection.
Our first result is a decomposition for the optimal token level selection rule that establishes a connection to the importance weighted sampling in Definition 1. The proof is in Appendix A.
Theorem 1 (Optimal Acceptance Probability and Canonical Architecture)
@@
In addition, if $\{\beta^{\star}_{y}(x_{1:K})\}$ denotes the parameters that achieve the maximum in (2), then $P^{\star}(\mathrm{acc})$ can be attained by a two step canonical architecture: in the first step, given the list of input tokens $\{x_{1},\ldots,x_{K}\}$, we apply Importance Weighted Sampling in Definition 1 with parameters $\beta^{\star}_{y}(x_{1},\ldots,x_{K})$ to output an intermediate token $y\in\{x_{1},\ldots,x_{K}\}$; in the second step we apply a single-draft speculative sampling scheme (Chen et al., 2023; Leviathan et al., 2023) on the selected token $y$ to generate the final output token.
@@
Figure 1 illustrates the proposed architecture Theorem 1, where the first step involves importance weighted sampling to output an intermediate token and the second step involves speculative sampling. This approach requires computing the optimal $\beta^{\star}_{y}(x_{1:K})$. In practice one can use sub-optimal choices that are faster to compute, as will be discussed in the sequel.
@@
Although our theoretical development focuses on the case of identical drafts, we can also easily extend Theorem 1 when the distributions of the $K$ tokens are not identical i.e., ${\mathcal{S}}\sim\prod_{i=1}^{K}p_{i}(X_{i})$. This is discussed in Section A.1 in the supplementary material.
@@
Note that the acceptance probability can equal $1$ even when $p(\cdot)$ and $q(\cdot)$ are not identical. Thus when the distribution of the draft model is close to the target model but not equal the acceptance probability can equal $1$. This is in contrast to the case of $K=1$, where it is known that the acceptance probability can only equal $1$ when $p(\cdot)$ and $q(\cdot)$ are identical distributions (Sun et al., 2024b). Furthermore to the best of our knowledge, previously proposed schemes for the multi-draft setting, such as SpecTr (Sun et al., 2024b) and SpecInfer (Miao et al., 2024) based on modified rejection sampling also require $p(\cdot)=q(\cdot)$ for the acceptance probability to be $1$. Theorem 1 is interesting in the context of our two-step architecture in Fig. 1. In this case, the output of importance weighted sampling block $Y$ matches the target distribution $q(\cdot)$ and the second step involving speculative sampling is not needed.
@@
The proof of Theorem 2 in Appendix B involves analyzing the output distribution $p_{I}(\cdot)$ of the Importance Weighted Sampling Scheme in Theorem 1 and demonstrating that a feasible choice of $\beta_{y}(x_{1},x_{2})$ exists and sets $p_{I}(\cdot)=q(\cdot)$ when the condition (5) is satisfied. The proof is based on the Fourier-Motzkin (FM) elimination technique (Ziegler, 2012). However a direct application of such a technique to satisfy the constraints $q(i)=p_{I}(i)$ for each $i\in\Omega$ becomes intractable. Our key idea is to demonstrate that instead considering a relaxation of the form $q(i)\geq p_{I}(i)$ leads to the same solution as the equality constraints and is amenable to analysis using Fourier-Motzkin elimination. We explain this further with an example involving $\Omega=\{1,2,3\}$ in Appendix B.
The core technical challenge in the proof of Theorem 2 is in determining whether a system of linear equations has a non-negative solution. Such problems been studied previously in the literature, with Chernikova (1964); Dines (1926) providing an algorithm. Such considerations lead to a geometric viewpoint involving polyhedral cones which we discuss in Appendix C. We explain how the double-description method (Fukuda and Prodon, 1995) for finding dual representations of polyhedral cones can be used to numerically verify the necessary and sufficient condition for the acceptance probability to equal $1$. In fact this approach was used to verify analogous conditions to Theorem 2 for up to $K=6$ drafts and all alphabets of size $|\Omega|\leq 14$, although we only provide an analytical proof of the condition for $K=2$ drafts in this paper. The reason for this is that our key step i.e., Lemma 2 in the Appendix that makes use of Fourier-Motzkin elimination, does not easily generalize to the case of $K>2$ drafts.
@@
To the best of our knowledge the result in Theorem 3 was not known before. Upper bounds on $P^{\star}(\mathrm{acc})$ are presented in Sun et al. (2024b), which are not necessarily tight. In contrast  (6) provides an exact expression for the acceptance probability for the case of $K=2$ drafts when $X_{1}$ and $X_{2}$ are independently sampled from $p(\cdot)$. The result in Theorem 3 implies the result in Theorem 2. In particular, note that if
$\sum_{s\in{\mathcal{S}}}q(s)\geq\left(\sum_{s\in{\mathcal{S}}}p(s)\right)^{2}$ (7)
it follows that:
$\sum_{s\in{\mathcal{S}}}q(s)+\left(\sum_{s\in{\mathcal{S}}^{c}}p(s)\right)^{2}+2\left(\sum_{s\in{\mathcal{S}}}p(s)\right)\left(\sum_{s\in{\mathcal{S}}^{c}}p(s)\right)\geq\left(\sum_{s\in{\mathcal{S}}}p(s)+\sum_{s\in{\mathcal{S}}^{c}}p(s)\right)^{2}=1,$ (8)
with equality when ${\mathcal{S}}=\Omega$. The proof of Theorem 3, presented in Appendix D, applies the Fourier-Motzkin elimination to the linear program presented in Theorem 1 to characterize an analytical solution in the case of $K=2$ drafts. The proof builds upon the proof of Theorem 2 but requires elimination of additional variables.

### v2

# Multi-Draft Speculative Sampling: Canonical Decomposition and Theoretical Limits

@@ Abstract
We consider multi-draft speculative sampling, where the proposal sequences are sampled independently from different draft models. At each step, a token-level draft selection scheme takes a list of valid tokens as input and produces an output token whose distribution matches that of the target model. Previous works have demonstrated that the optimal scheme (which maximizes the probability of accepting one of the input tokens) can be cast as a solution to a linear program. In this work we show that the optimal scheme can be decomposed into a two-step solution: in the first step an importance sampling (IS) type scheme is used to select one intermediate token; in the second step (single-draft) speculative sampling is applied to generate the output token. For the case of two identical draft models we further 1) establish a necessary and sufficient condition on the distributions of the target and draft models for the acceptance probability to equal one and 2) provide an explicit expression for the optimal acceptance probability. Our theoretical analysis also motives a new class of token-level selection schemes based on weighted importance sampling. Our experimental results demonstrate consistent improvements in the achievable block efficiency and token rates over baseline schemes in a number of scenarios.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has revolutionized the field of natural language processing and deep learning. One of the key factors contributing to the success story of transformers, as opposed to prior recurrent-based architectures (Hochreiter and Schmidhuber, 1997, Chung et al., 2014), is their inherent train-time parallelization due to the attention mechanism. This allows for massive scaling and lead to the development of state-of-the-art Large Language Models (LLMs) (Touvron et al., 2023, Achiam et al., 2023, Brown et al., 2020, Chowdhery et al., 2023) which have demonstrated remarkable performance across a wide range of tasks. Despite their parallelizable training, LLM inference is sequential, owing to their auto-regressive nature. This limits their text-generation to one token per one forward pass, which is known to be memory-bound (Shazeer, 2019).
To alleviate the memory-bound nature of auto-regressive decoding of LLMs, speculative decoding (Chen et al., 2023, Leviathan et al., 2023) leverages an arbitrary smaller language model (draft model) that generates multiple candidate tokens in an auto-regressive manner. The LLM (target model) is then used to score all the tokens in the draft in parallel, and the draft tokens are verified through a sequence of token-level rejection sampling which guarantees that the final sequence follows the same distribution as that of the target model. In order for speculative decoding to be beneficial, the combined cost of auto-regressively sampling from the draft model and parallel verification via the target model should be smaller than auto-regressively sampling from the target model. Intuitively, this requires that the draft model distribution resembles that of the target model, which can be measured via the acceptance rate of the speculative decoding process, i.e., the rate at which we accept/reject draft tokens.
A large number of works on speculative decoding (Sun et al., 2024b, Jeon et al., 2024, Miao et al., 2024, Sun et al., 2024a) have emerged recently in an effort to further improve decoding efficiency. The authors in (Sun et al., 2024b) propose SpecTr, a multi-draft extension where the draft model generates $K$ candidate token sequences (which could be sampled in a batch) for each time-step (as opposed to one). The authors consider a token-level selection scheme with the objective of maximizing the probability of accepting some token in the set of available tokens. They demonstrate that this problem can be cast into the framework of optimal transport and solved using a linear program. However due to complexity reasons, the authors instead propose a modified sequential rejection sampling scheme.
@@ 1.1 Main Contributions
  • We revisit the optimal transport framework introduced in  (Sun et al., 2024b) and introduce a canonical decomposition — we demonstrate that the optimal acceptance probability can be achieved by a two-step scheme: the first step involves selecting a token from the input set using a type of importance sampling; the second step involves speculative sampling using the selected token and the target distribution.
@@ 2 Background and Related Work
Auto-regressive sampling from LLMs is inherently sequential and memory-bound (Shazeer, 2019). Several approaches have been proposed in the literature to accelerate LLM inference (Shazeer, 2019, Jaszczur et al., 2021, Frantar et al., 2022, Frantar and Alistarh, 2023, Stern et al., 2018, Chen et al., 2023, Leviathan et al., 2023, Jeon et al., 2024, Sun et al., 2024b, Miao et al., 2024). Model compression techniques, such as quantization (Frantar et al., 2022, Bondarenko et al., 2024) and sparsification (Jaszczur et al., 2021, Frantar and Alistarh, 2023) have been shown to reduce the overall complexity of LLMs at the expense of some degradation in decoding quality.
@@
For lossless LLM inference acceleration, speculative decoding (Chen et al., 2023, Leviathan et al., 2023, Stern et al., 2018) has emerged as a promising and orthogonal alternative. Earlier works on greedy decoding can draft and predict multiple tokens by augmenting the base LLM (Stern et al., 2018) or aggressive decoding (Ge et al., 2022). However, LLM text-generation often requires sampling with non-zero temperature from the generated logits. To that end, speculative decoding (Chen et al., 2023, Leviathan et al., 2023) was proposed. In speculative decoding, auto-regressive sampling is delegated to a smaller language model (draft model) that generates multiple candidate tokens. The LLM (target model) is then used to score all the tokens in the draft in parallel, and the draft tokens are verified through a sequence of token-level rejection sampling. Speculative decoding guarantees that the final sequence follows the same distribution as that of the target model. The performance of speculative methods highly depends on the choice of the draft model. Zhou et al. (2023) use knowledge distillation (Hinton et al., 2015) to better align the draft and target models which results in higher token acceptance rates.
@@
More recently, the works of Sun et al. (2024b), Miao et al. (2024), Jeon et al. (2024) extend speculative decoding to the multi-draft setting where the draft model(s) generate multiple token sequences per time-step. Specifically, Sun et al. (2024b) formulate the token-level draft selection problem as a discrete optimal transport problem with membership cost and propose SpecTr: a new decoding algorithm that allows for multiple candidates for each token in the draft. A related setting is also studied in Miao et al. (2024), Jeon et al. (2024) where the authors consider a token tree based construction for improving the draft sequences as well as a token-level selection method different form Sun et al. (2024b). Instead of using a dedicated draft model, Cai et al. (2024) propose augmenting the target model with extra decoding heads that can concurrently draft multiple tokens. The extra heads are fine-tuned using parameter-efficient methods, and can be added to any pre-trained target model. Orthogonally, Sun et al. (2024a) study block-level verification in the single-draft setting as a block-level optimal transport problem. They propose a computationally-efficient algorithm that optimally solves the block-level transport problem, and report speedups over prior token-level verification (Leviathan et al., 2023).
@@ 3 Token-Level Optimal Draft Selection: Theoretical Analysis
We focus on token-level optimal draft selection framework introduced in Sun et al. (2024b). For sake of completeness we review the speculative sampling schemes in Leviathan et al. (2023), Chen et al. (2023) in Appendix A. We assume that $\Omega=\{1,2,\ldots,n\}$ denotes the vocabulary of tokens and at a given step, say $t$, ${\mathcal{S}}=\{X_{1},\ldots,X_{K}\}$, denotes the $K$ valid tokens under consideration. Each of these tokens is generated in an i.i.d. fashion from a distribution $p(\cdot)$ determined by the underlying draft modelWe also consider the case when the tokens are generated from different distributions. See Remark 1 as well as the experimental results in Section 5. and the context sequence $u^{t}\in\Omega^{t}$ i.e, for each $y\in\Omega$, we have $p(y)={\mathcal{M}}_{s}(y|u^{t})$, where ${\mathcal{M}}_{s}$ denotes the distribution generated by the small (draft) model. In a similar fashion we let $q(\cdot)$ be the distribution over $\Omega$ associated with the large model i.e., $q(y)={\mathcal{M}}_{b}(y|u^{t})$ where ${\mathcal{M}}_{b}$ denotes the distribution generated by the large model. Note that we do not explicitly indicate the sequence $u^{t}$ when discussing $p(\cdot)$ and $q(\cdot)$, as it is fixed and common to both models throughout our analysis.
@@
Problem 1 was studied in Sun et al. (2024b) and shown to be an instance of optimal transport, which can be cast as a linear program. The authors used this framework to establish the optimality of speculative sampling (Chen et al., 2023, Leviathan et al., 2023) in the case of a single draft i.e., $K=1$. For $K>1$ the authors established an information theoretic upper bond on $P^{\star}(\mathrm{acc})$. In this work, we revisit Problem 1 and develop new insights into the structure of the optimal solution. In fact, we establish that the optimal solution in the case of multiple drafts has a natural connection to importance sampling (Tokdar and Kass, 2010). For the case of $K=2$ drafts we exactly characterize $P^{\star}(\mathrm{acc})$ and state necessary and sufficient conditions on $p(\cdot)$ and $q(\cdot)$ for $P^{\star}(\mathrm{acc})$ to equal $1$.
@@
Note that instead of considering the probability over the value of the selected token in (1), one can instead consider the probability of selecting an index $i$ between $\{1,\,\ldots,K\}$ i.e., $\Pr(I=i|X_{1:K}=x_{1:K})$. Such a distribution maps to (1) by simply summing over all indices where $x_{i}=y$. We note that the form in (1) will be more convenient in the sequel. Also note that the classical importance sampling scheme (Tokdar and Kass, 2010) corresponds to the case where $\Pr(I=i|X_{1:K}=x_{1:K})\propto q(x_{i})/p(x_{i})$. However the family of schemes in Definition 1 is not restricted to such a choice and we treat $\beta_{y}(x_{1},\ldots,x_{K})$ as free parameters that can be optimized. While an importance weighted sampling scheme may not lead be a valid TLSR, as explained next, it is a key building block in a canonical decomposition for token selection.
Our first result is a decomposition for the optimal token level selection rule that establishes a connection to the importance weighted sampling in Definition 1. The proof is in Appendix B.
Theorem 1 (Optimal Acceptance Probability and Canonical Decomposition)
@@
In addition, if $\{\beta^{\star}_{y}(x_{1:K})\}$ denotes the parameters that achieve the maximum in (2), then $P^{\star}(\mathrm{acc})$ can be attained by a two step canonical decomposition: in the first step, given the list of input tokens $\{x_{1},\ldots,x_{K}\}$, we apply Importance Weighted Sampling in Definition 1 with parameters $\beta^{\star}_{y}(x_{1},\ldots,x_{K})$ to output an intermediate token $y\in\{x_{1},\ldots,x_{K}\}$; in the second step we apply a single-draft speculative sampling scheme (Chen et al., 2023, Leviathan et al., 2023) on the selected token $y$ to generate the final output token.
@@
Figure 1 illustrates the proposed system in Theorem 1, where the first step involves importance weighted sampling to output an intermediate token and the second step involves speculative sampling. This approach requires computing the optimal $\beta^{\star}_{y}(x_{1:K})$. In practice one can use sub-optimal choices that are faster to compute, as will be discussed in the sequel.
@@
Although our theoretical development focuses on the case of identical drafts, our result in Theorem 1 naturally extends when the the $K$ tokens are instead sampled from a joint distribution i.e., $p(x_{1},\ldots x_{K})$. This is discussed in Section B.1 in the supplementary material. In particular Theorem 1 naturally extends to the setting when the $K$ tokens are sampled independently from different distributions ${\mathcal{S}}\sim\prod_{i=1}^{K}p_{i}(X_{i})$ as in Miao et al. (2024) as well as to the case when the tokens are sampled without replacement (Jeon et al., 2024).
@@
Note that the acceptance probability can equal $1$ even when $p(\cdot)$ and $q(\cdot)$ are not identical. Thus when the distribution of the draft model is close to the target model but not equal the acceptance probability can equal $1$. This is in contrast to the case of $K=1$, where it is known that the acceptance probability can only equal $1$ when $p(\cdot)$ and $q(\cdot)$ are identical distributions (Sun et al., 2024b). Furthermore to the best of our knowledge, previously proposed schemes for the multi-draft setting, such as SpecTr (Sun et al., 2024b) and SpecInfer (Miao et al., 2024) based on modified rejection sampling also require $p(\cdot)=q(\cdot)$ for the acceptance probability to be $1$. Theorem 1 is interesting in the context of our two-step system in Fig. 1. In this case, the output of importance weighted sampling block $Y$ matches the target distribution $q(\cdot)$ and the second step involving speculative sampling is not needed.
@@
The proof of Theorem 2 in Appendix C involves analyzing the output distribution $p_{I}(\cdot)$ of the Importance Weighted Sampling Scheme in Theorem 1 and demonstrating that a feasible choice of $\beta_{y}(x_{1},x_{2})$ exists and sets $p_{I}(\cdot)=q(\cdot)$ when the condition (5) is satisfied. The proof is based on the Fourier-Motzkin (FM) elimination technique (Ziegler, 2012). However a direct application of such a technique to satisfy the constraints $q(i)=p_{I}(i)$ for each $i\in\Omega$ becomes intractable. Our key idea is to demonstrate that instead considering a relaxation of the form $q(i)\geq p_{I}(i)$ leads to the same solution as the equality constraints and is amenable to analysis using Fourier-Motzkin elimination. We explain this further with an example involving $\Omega=\{1,2,3\}$ in Appendix C.
The core technical challenge in the proof of Theorem 2 is in determining whether a system of linear equations has a non-negative solution. Such problems have been studied previously in the literature, with Chernikova (1964), Dines (1926) providing an algorithm. Such considerations lead to a geometric viewpoint involving polyhedral cones which we discuss in Appendix D. We explain how the double-description method (Fukuda and Prodon, 1995) for finding dual representations of polyhedral cones can be used to numerically verify the necessary and sufficient condition for the acceptance probability to equal $1$. In fact this approach was used to verify analogous conditions to Theorem 2 for up to $K=6$ drafts and all alphabets of size $|\Omega|\leq 14$, although we only provide an analytical proof of the condition for $K=2$ drafts in this paper. The reason for this is that our key step i.e., Lemma 2 in the Appendix that makes use of Fourier-Motzkin elimination, does not easily generalize to the case of $K>2$ drafts.
@@
To the best of our knowledge the result in Theorem 3 was not known before. Upper bounds on $P^{\star}(\mathrm{acc})$ are presented in Sun et al. (2024b), which are not necessarily tight. In contrast  (6) provides an exact expression for the acceptance probability for the case of $K=2$ drafts when $X_{1}$ and $X_{2}$ are independently sampled from $p(\cdot)$. The proof of Theorem 3, presented in Appendix E, applies the Fourier-Motzkin elimination to the linear program presented in Theorem 1 to characterize an analytical solution in the case of $K=2$ drafts. The proof builds upon the proof of Theorem 2 but requires elimination of additional variables.